OpenStudy (anonymous):

PLEASE HELP ME :( prove limit of (x^3*y)/(sqrt(x^8+y^8) does not exist

OpenStudy (amistre64):

limit as x,y approach ? you essentially have to approach it from all directions and paths

OpenStudy (anonymous):

yes, I have tried from x=0, y=0, y=mx, etc. I cannot find an equation where it will not approach 0 as the limit

OpenStudy (anonymous):

it is the limit as (x,y) approach (0,0)

OpenStudy (amistre64):

\[\large\frac{x^3y}{\sqrt{x^8+y^8~}}\] hmm, can you do LHop on it?

OpenStudy (anonymous):

I'm not sure how that would work in this type of problem

OpenStudy (amistre64):

me either ... something to do with Fx Fy Fxy Fyx perhaps

OpenStudy (amistre64):

try a quadratic approach? y = m(x-a)^2+b

OpenStudy (amistre64):

of course the +b is prolly 0 for that ...

OpenStudy (amistre64):

the limit of y=x is 1/sqrt(2) the limit of y=2x is 2/sqrt(257) the limit is different depending on the path taken ...

OpenStudy (amistre64):

different limits means no limit can be defined

OpenStudy (amistre64):

let y = mx \[\large\frac{mx^4}{x^4\sqrt{1+m^8}}\] \[\lim_{x\to 0} \large\frac{m}{\sqrt{1+m^8}}\]is not the same for all m

OpenStudy (anonymous):

when I did the limit as x approaches 0 with y=mx I got 0, was that wrong?

OpenStudy (amistre64):

im going to go out on a limb and say .... yes, you were wrong :)

OpenStudy (amistre64):

the xs cancel out, leaving us a limit that differs from m to m

OpenStudy (anonymous):

how did you get the limit of y=x is 1/sqrt(2) the limit of y=2x is 2/sqrt(257), I am getting the limit with y=x as x approaches 0 to equal 1/sqrt(x^4) which, when x = 0 wouldn't give you that?

OpenStudy (amistre64):

for y = mx \[\frac{m}{\sqrt{1+m^8}}\] \[\frac{1}{\sqrt{1+1}}\] \[\frac{2}{\sqrt{1+2^8}}\] \[\frac{3}{\sqrt{1+3^8}}\] for any given direction of m, we get a different limit value

OpenStudy (anonymous):

how are you getting from mx^4/sqrt(x^8+mx^8) to that original part?

OpenStudy (anonymous):

mx^4/sqrt(x^8(1+m^8)) then what, you can factor out x^4??

OpenStudy (amistre64):

the most basic definition of a limit is that it must approach the same value regardless of direction ... otherwise it does not exist or cannot be calculated \[\large\frac{x^3y}{\sqrt{x^8+y^8~}}\] \[y=mx\] \[\large\frac{x^3(mx)}{\sqrt{x^8+(mx)^8~}}\] \[\large\frac{mx^4}{\sqrt{x^8+m^8x^8~}}\] \[\large\frac{mx^4}{\sqrt{x^8(1+m^8)~}}\] \[\large\frac{mx^4}{\sqrt{x^8}\sqrt{1+m^8~}}\] \[\large\frac{mx^4}{{x^{8/2}}\sqrt{1+m^8~}}\] \[\large\frac{m\cancel{x^4}}{\cancel{x^{4}}\sqrt{1+m^8~}}\]

OpenStudy (amistre64):

the limit is not the same for all values of m, therefore it DNE or cannot be determined

OpenStudy (anonymous):

THANK YOU so much!! I have been having trouble with this problem for the past week and it's due with a problem set tomorrow... I really appreciate your help!!